For the 2nd order system given Ts =7 sec and Tp= 3 sec find the value of % OS and location of poles.


1
Expert's answer
2021-08-16T02:19:52-0400

"T_s= \frac{4}{\xi w_n}\\
w_n= \frac{0.571}{\xi}\\
\xi= \frac{\pi}{\frac{0.571}{\xi} \sqrt{1-\xi^2}}\\
\sqrt{1- \xi^2}=(\frac{\pi}{3(0.571)}) \xi\\
1-4363 \xi^2\\
\xi= 0.479\\
\therefore w_n= \frac{4}{7*0.479}= 1.193 rad/sec\\
s^2+(2*0.479*1.193)s+1.193)s+1.193^2=0\\
s= -\frac{1.143±\sqrt{1.143^2-4*1.423}}{2}\\
s=-0.542±j0.903"